OpenStudy (anonymous):

the distance of a moving point from the line x=3 is twice its distance from (-1,-2). FInd the equation of its locus.

OpenStudy (anonymous):

let P=(x,y) d1^2 -> use distance formula (x-3=0,(x,y))=|x-3|/sqrt(1)=(x-3)^2 d2^2 -> use distance formula between two points d1^2 = 4*d2^2

OpenStudy (anonymous):

can you please give an illustration?

OpenStudy (anonymous):

and i really don't get it... -_- i'm sorry.. can you please elaborate

OpenStudy (anonymous):



OpenStudy (anonymous):

d1 is distance of (x,y) from the line x=3 d2 is distance between (x,y) and (-1,-2)

OpenStudy (anonymous):

how will i find for the equation of the locus?

OpenStudy (anonymous):

find d1, d2 it is given that d1=2*d2 you'd get an equation in terms of x and y do you know what a locus is?

OpenStudy (anonymous):

it is a set of points

OpenStudy (anonymous):

yes but satisfying certain conditions.

OpenStudy (anonymous):

yes.. ok.. i'll try solving that

OpenStudy (anonymous):

so, all possible (x,y) that satisfy one or many equations. that equation is what the question is asking you to find,

OpenStudy (anonymous):

d1^2 -> use distance formula (x-3=0,(x,y))=|x-3|/sqrt(1)=(x-3)^2 how did you do this?

OpenStudy (anonymous):

distance of a point from a line \[ l\implies ax+by+c=0\quad (x_0,y_0)\\ d_1=\frac{|ax_0+by_0+c|}{\sqrt{a^2-b^2}} \]

OpenStudy (anonymous):

ok ok.. got it.. but do i really have to square both d to find the locus?

OpenStudy (anonymous):

makes life easy.. d2 is a square-root remember?

OpenStudy (anonymous):

what i got was 3x^2 + 14x + 4y^2 + 16y + 25 = 0.. is this correct?

OpenStudy (anonymous):

3x^2 + 14x + 4y^2 + 16y + 7 = 0..

OpenStudy (anonymous):

hope so. ;)

OpenStudy (anonymous):

unless you did some mistake with the numbers.

OpenStudy (anonymous):

hahaha.. ok .. i'll write what i did so you can see which part i made the mistake..

OpenStudy (anonymous):

it's so hard to write it down.. :(

OpenStudy (anonymous):

i know.. but its cool. should look something like that in fact, it is a circle.

OpenStudy (anonymous):

or an ellipse?.. the co-efficients of x^2 and y^2 are different.. so, an ellipse

OpenStudy (anonymous):

3x^2 - 2x + 4y^2 + 8y - 3 = 0.. that's what i got.. is it correct?

OpenStudy (anonymous):

3x^2 + 2x + 4y^2 + 8y - 3 = 0

OpenStudy (anonymous):

yes. you can check. put some x, like "2", find y check distances

OpenStudy (anonymous):

thank you so much! :))